rewrite equation so the y is a function of x
3x + 18y = 36

Answer by oscargut(2103)


3x + 18y = 36
18y = 36-3x
y = (36-3x)/18
Answer: y = 2 -(1/6)x
If you have more questions i can see them if you contact me at: mthman@gmail.com
Thanks